# Radio-frequency assisted switching in perpendicular magnetic tunnel junctions

### v1

@@ Discussion
Figure 3 reveals that RF-assisted switching enhancement is observed for both positive and negative delays $\tau$ between the RF and DC pulses. The enhancement for $\tau<0$ can be expected because the overlap between the RF and DC pulses can result in the net applied voltage exceeding $V_{\mathrm{DC}}$ in an interval of the combined RF+DC waveform. However, the RF-enhanced switching in the $\tau\geq 0$ regime is surprising. This latter regime is also important from a practical point of view because the combined RF+DC waveform voltage never exceeds $V_{\mathrm{DC}}$, leading to enhanced switching without extra stress on the junction. From another angle, introduction of an RF pulse allows one to reduce the DC pulse duration $t_{\mathrm{DC}}$ for a targeted switching probability compared to the DC-only protocol. Since longer DC write pulses are known to stress the MTJ barrier oxide and reduce device endurance [46, 47], the RF+DC switching protocol in the $\tau\geq 0$ regime studied here is expected to enhance STT-MRAM endurance.
The maximum RF pulse amplitude used in this study, $V_{\mathrm{RF}}=0.195$ V, is much smaller than the DC write pulse amplitude $\sim 1$ V. Therefore, in the $\tau\geq 0$ regime, this small-amplitude RF assist does not diminish device endurance and only marginally increases the write energy while increasing the switching probability. This makes RF assist at $\tau=0$ a viable strategy for an improved STT-MRAM write protocol.
@@
In principle, ohmic heating induced by the RF pulse could contribute to the enhancement of switching probability $\Delta\mathcal{P}$. However, because ohmic heating is frequency independent, it cannot account for the strong frequency dependence seen in Fig. 5. Circuit calibrations used in our experiments ensured uniform RF pulse power across all frequencies. If heating were the dominant mechanism, the observed switching enhancement would be comparable across all frequencies at constant power. The strong frequency dependence of the RF assist demonstrated in Fig. 5 therefore eliminates ohmic heating of the MTJ by the RF pulse as a possible mechanism of the observed RF-induced switching enhancement [48].
@@
The theoretically predicted $\sim 1$ ns time scale is reasonable because it is similar to the characteristic magnetic energy relaxation time of the FL $\tau_{E}=1/\left(4\pi\alpha f_{\mathrm{FMR}}\right)=0.66$ ns, where $\alpha=0.0097$ and $f_{\mathrm{FMR}}=12$ GHz as determined by spin torque FMR measurements (Supplementary Note 2). For $\tau\gg\tau_{E}$, all magnetic energy added to the FL magnetization by the RF pulse is expected to dissipate away, and thus magnetization dynamics driven by the DC pulse is expected to be unaffected by the RF pulse. The observation of RF assist effects for $\tau\gg\tau_{E}$ suggests that, in addition to FMR dynamics, the FL also supports significantly slower magnetic dynamics.
Signatures of such slow magnetic dynamics have been previously observed in a different experiment [49] and were attributed to regions at the FL/MgO interface with reduced magnetization and weak exchange coupling. Our data in Fig. 4 provide additional support for this hypothesis of slow dynamics in the FL of a p-MTJ. Moreover, the greater enhancement of switching probability $\Delta\mathcal{P}$ at lower RF frequencies as shown in Fig. 5 suggests these dynamics are more efficiently excited at lower frequencies. This effect may be rooted in the structural, chemical, and magnetic inhomogeneity of the ultrathin CoFeB FL found in MTJ devices [50]. Electrical transport, FMR and magnetometry studies on ultrathin films and MTJ nanopillars based on Ta/CoFeB/MgO multilayers suggest that iron-based magnetic oxides form at the CoFeB/MgO interface [50, 51]. These slow magnetic dynamics attributed to the FL inhomogeneity could contribute to the enhancement of switching probability $\Delta\mathcal{P}$ for $\tau\geq 0$ but are beyond the scope of the theoretical framework of this paper.
@@
In this work, we only made measurements of RF-assisted switching for random phase of the RF signal with respect to the onset of the DC write pulse. Our theoretical results in Fig. 6a,b suggest a route toward efficient RF-assisted switching in the regime of small overlap between the RF and DC pulses $\tau<0$ when the phase of the RF signal is fixed. Indeed, the predicted switching probability in Fig. 6b shows switching probability enhancement nearing $\sim 50\%$ for 3 ns overlap of the RF and DC pulses. This is an order of magnitude larger than for the case of random RF pulse phase in Fig. 6d. While the RF-DC overlap regime may decrease the MTJ endurance, this decrease may be relatively small in some regimes of STT-MRAM operation, and the greatly increased switching probability may prove to be more important. This observation calls for future studies of the RF-assisted switching with fixed phase of the RF pulse.
@@ Methods
RF-assisted switching probability measurements Current-induced switching of the MTJ FL is observed by monitoring the change in the MTJ resistance in a circuit where the MTJ forms a voltage divider with a fixed resistor. We bias the voltage divider by a low-level DC voltage and use a 16-bit 2 MHz data acquisition board (DAQ) to measure time-dependent voltage across the MTJ. To switch the FL, we apply a DC voltage pulse preceded by an RF voltage pulse to the MTJ. Details of the circuit and the waveform protocol are given in Supplementary Note 1. Following the RF+DC write waveform, we use DAQ to measure the resultant resistance state of the MTJ, which registers either switching or non-switching. A lower-amplitude, longer-duration DC reset pulse of the opposite polarity to the DC write pulse is then used to return the device to its initial state, and the DAQ is used to confirm the successful reset.
The RF pulse is generated by using a broadband microwave mixer connected to a microwave generator controlling the RF frequency and a square pulse generator controlling the output RF pulse duration. The amplitude of the RF pulse is controlled by both generators. All three pulse generators used in the measurements are synchronized with each other and the DAQ as described in Supplementary Note 1. The microwave generator operates in continuous wave mode, resulting in a random phase $\phi$ of the RF pulse. The delay $\tau$ between the RF and DC pulses is controlled through a combination of the cable length and a programmable internal delay of the DC write pulse generator. The switching probability $\mathcal{P}$ is determined as a ratio of the number of successful switching events to the total number of switching attempts. The voltage levels of all RF+DC waveforms used in the measurements are characterized using a high-bandwidth real-time oscilloscope.
@@
All data generated or analyzed during this study are included in this published article and are available from the corresponding author on reasonable request.
@@ Acknowledgements
This work was primarily supported by Western Digital Technologies. Funding from the National Science Foundation via awards ECCS-2213690 and DMREF-2324203 as well as from the University of California National Laboratory Fees Research Program is also acknowledged. The authors thank the Eddleman Quantum Institute (EQI) for partial support of this work. S.P gratefully acknowledges financial support for this publication by the Fulbright Visiting Scholar Program, which is sponsored by the U.S. Department of State and the Italian Ministry of Foreign Affairs and administered by the U.S. - Italy Fulbright Commission. Its contents are solely the responsibility of the author and do not necessarily represent the official views of the Fulbright Program, the Government of the United States, or the U.S. - Italy Fulbright Commission.

### v2

@@ Discussion
Figure 3 reveals that RF-assisted switching enhancement is observed for both positive and negative delays $\tau$ between the RF and DC pulses. The enhancement for $\tau<0$ can be expected because the overlap between the RF and DC pulses can result in the net applied voltage exceeding $V_{\mathrm{DC}}$ in an interval of the combined RF+DC waveform. However, the RF-enhanced switching in the $\tau\geq 0$ regime is surprising. This latter regime is also important from a practical point of view because the combined RF+DC waveform voltage never exceeds $V_{\mathrm{DC}}$, leading to enhanced switching without extra stress on the junction. From another angle, introduction of an RF pulse allows one to reduce the DC pulse duration $t_{\mathrm{DC}}$ for a targeted switching probability compared to the DC-only protocol. Since longer DC write pulses are known to stress the MTJ barrier oxide and reduce device endurance [1, 19, 50], the RF+DC switching protocol in the $\tau\geq 0$ regime studied here is expected to enhance STT-MRAM endurance. Prior studies of STT-MRAM reported a nearly order-of-magnitude improvement in the MgO write cycles before the dielectric breakdown when $t_{\mathrm{DC}}$ was reduced from $10$ ns to $5$ ns [50].
The maximum RF pulse amplitude used in this study, $V_{\mathrm{RF}}=0.195$ V, is much smaller than the DC write pulse amplitude $\sim 1$ V. Therefore, in the $\tau\geq 0$ regime, this small-amplitude RF assist does not diminish device endurance and only marginally increases the write energy while increasing the switching probability. This makes RF assist at $\tau=0$ a viable strategy for an improved STT-MRAM write protocol. For example, addition of a small-amplitude RF pulse at $\tau=0$ in Fig. 4a, leads to a modest 17 % increase of the write energy while giving a 13 % increase in the switching probability without affecting the device endurance. Our calculations of the energy per write (Supplementary Note 1 and Figure S2) identify regimes in which the RF+DC and DC protocols have equal write energies, while $t_{\mathrm{DC}}$ is shorter in the RF+DC protocol. Figure S2 shows an example where the addition of a small-amplitude RF pulse reduces $t_{\mathrm{DC}}$ while maintaining the same total write energy for $\tau=-7$ ns. A comprehensive follow-up study of STT switching as a function of RF and DC pulse parameters is required to determine whether the RF+DC protocol can achieve higher energy efficiency than the DC-only write protocol. The present study focuses on regimes $\tau\geq 0$ where the RF+DC protocol improves the switching probability without compromising STT-MRAM endurance.
@@
In principle, ohmic heating induced by the RF pulse can contribute to the enhancement of switching probability $\Delta\mathcal{P}$. However, because time-averaged ohmic heating is frequency independent, MTJ heating averaged over the RF pulse duration cannot account for the strong frequency dependence seen in Fig. 5. Circuit calibrations used in our experiments ensured uniform RF pulse power across all frequencies (See Supplementary Note 1 and Table S1). If time-averaged MTJ temperature increase induced by the RF pulse were the dominant mechanism, the observed switching enhancement would be comparable across all frequencies at constant power. The strong frequency dependence of the RF assist demonstrated in Fig. 5 therefore eliminates time-averaged ohmic heating of the MTJ by the RF pulse as the primary mechanism of the observed RF-induced switching enhancement [34].
@@
The theoretically predicted $\sim 1$ ns time scale is reasonable because it is similar to the characteristic magnetic energy relaxation time of the FL $\tau_{E}=1/\left(4\pi\alpha f_{\mathrm{FMR}}\right)=0.66$ ns, where $\alpha=0.0097$ and $f_{\mathrm{FMR}}=12$ GHz as determined by spin torque FMR measurements (Supplementary Note 2 and Figure S3). For $\tau\gg\tau_{E}$, all magnetic energy added to the FL magnetization by the RF pulse is expected to dissipate away, and thus magnetization dynamics driven by the DC pulse is expected to be unaffected by the RF pulse. The observation of RF assist effects for $\tau\gg\tau_{E}$ suggests that, in addition to FMR dynamics, the FL also supports significantly slower magnetic dynamics.
Signatures of slow magnetic dynamics have been previously observed in a different experiment [39] and were attributed to regions at the FL/MgO interface with reduced magnetization and weak exchange coupling. Our data in Fig. 4 provide additional support for this hypothesis of slow dynamics in the FL of a p-MTJ. Moreover, the greater enhancement of switching probability $\Delta\mathcal{P}$ at lower RF frequencies as shown in Fig. 5 suggests these dynamics are more efficiently excited at lower frequencies. This effect may be rooted in the structural, chemical, and magnetic inhomogeneity of the ultrathin CoFeB FL found in MTJ devices [5]. Electrical transport, FMR and magnetometry studies on ultrathin films and MTJ nanopillars based on Ta/CoFeB/MgO multilayers suggest that iron-based magnetic oxides form at the CoFeB/MgO interface [5, 4]. These slow magnetic dynamics attributed to the FL inhomogeneity could contribute to the enhancement of switching probability $\Delta\mathcal{P}$ for $\tau\geq 0$ but are beyond the scope of the theoretical framework of this paper.
@@
A potential mechanism of switching probability enhancement not included in the present theoretical description is the dynamics of the FL temperature and its coupling with the magnetization switching dynamics. An RF current applied to the MTJ produces a temperature increase that has a time-independent component that also does not depend on $f_{\mathrm{RF}}$ and an oscillatory component at the frequency $2f_{\mathrm{RF}}$ with a frequency-dependent amplitude approximately given by $T_{0}/\sqrt{1+(4\pi\tau_{\mathrm{\theta}}f_{\mathrm{RF}})^{2}}$, where $\tau_{\mathrm{\theta}}$ is the thermal time constant of the system and $T_{0}$ is the temperature oscillation amplitude in the low frequency limit [44, 36]. If we assume that the frequency dependence of the RF assisted switching in Fig. 5 is linked to these time-dependent temperature oscillations, we can fit the data in Fig. 5 to $\Delta\mathcal{P}_{0}/\sqrt{1+(4\pi\tau_{\mathrm{\theta}}f_{\mathrm{RF}})^{2}}$ in order to extract $\tau_{\mathrm{\theta}}$, where $\Delta\mathcal{P}_{0}$ is the switching probability enhancement in the low frequency limit. Such a fitting procedure gives $\tau_{\mathrm{\theta}}=0.18\pm 0.01$ ns.
Such a small thermal time constant is an order of magnitude below previously reported values of $\tau_{\mathrm{\theta}}$ for nanoscale MTJs [9]. In addition, $\tau\gg\tau_{\theta}$ is inconsistent with the data in Fig. 4, which shows that the effects of the RF pulse on the switching dynamics persists for nanoseconds after the end of the RF pulse. Note that the RF pulses of the magnitude and duration used in this study do not produce any measurable MTJ switching on their own without a follow-up DC pulse.
We, therefore, conclude that while a contribution of Ohmic heating to the observed RF assist cannot be completely excluded, it cannot be a simple process governed by a single time constant. Investigating a more complex mechanism would require careful measurements of the thermal time constants together with the development of a theory describing coupled thermal and magnetization dynamics. This analysis is beyond the scope of the present work and will be addressed in future studies.
@@ Methods
RF-assisted switching probability measurements Current-induced switching of the MTJ FL is observed by monitoring the change in the MTJ resistance in a circuit where the MTJ forms a voltage divider with a fixed resistor. We bias the voltage divider by a low-level DC voltage and use a 16-bit 2 MHz data acquisition board (DAQ) to measure time-dependent voltage across the MTJ. To switch the FL, we apply a DC voltage pulse preceded by an RF voltage pulse to the MTJ. Details of the circuit and the waveform protocol are given in Supplementary Note 1. Following the RF+DC write waveform, we use DAQ to measure the resultant resistance state of the MTJ, which registers either switching or non-switching. A lower-amplitude, longer-duration DC reset pulse of the opposite polarity to the DC write pulse is then used to return the device to its initial state, and the DAQ is used to confirm the successful reset. The RF pulse is generated by using a broadband microwave mixer connected to a microwave generator controlling the RF frequency and a square pulse generator controlling the output RF pulse duration. The amplitude of the RF pulse is controlled by both generators. All three pulse generators used in the measurements are synchronized with each other and the DAQ as described in Supplementary Note 1. The microwave generator operates in continuous wave mode, resulting in a random phase $\phi$ of the RF pulse. The delay $\tau$ between the RF and DC pulses is controlled through a combination of the cable length and a programmable internal delay of the DC write pulse generator. The switching probability $\mathcal{P}$ is determined as a ratio of the number of successful switching events to the total number of switching attempts. The voltage levels of all RF+DC waveforms used in the measurements are characterized using a high-bandwidth real-time oscilloscope.
@@
All data generated or analyzed during this study are included in this published article and are available from the corresponding author upon reasonable request.
@@ Acknowledgements
I.N.K. and P.M.B. thank Jordan Katine for stimulating discussions.
This work was primarily supported by Western Digital Technologies (grant number: not applicable). Funding from the National Science Foundation via awards ECCS-2213690 and DMREF-2324203 as well as from the University of California National Laboratory Fees Research Program (grant number: not applicable) is also acknowledged. The authors thank the Eddleman Quantum Institute (EQI) for partial support of this work (grant number: not applicable). S.P gratefully acknowledges financial support for this publication by the Fulbright Visiting Scholar Program, which is sponsored by the U.S. Department of State and the Italian Ministry of Foreign Affairs and administered by the U.S. - Italy Fulbright Commission (grant number: not applicable). The contents of this publication are solely the responsibility of the author and do not necessarily represent the official views of the Fulbright Program, the Government of the United States, or the U.S. - Italy Fulbright Commission.